OpenStudy (anonymous):

How do I find the integral of 3cos(x)/sin(x) ?

OpenStudy (isaiah.feynman):

Please write it with the equation editor so I will be clear.

OpenStudy (anonymous):

\[\int\limits_{\frac{ \pi }{ 6 }}^{\frac{ \pi }{ 2 }}\frac{ 3\cos(x) }{ \sin(x) } dx\]

OpenStudy (isaiah.feynman):

Looks fairly easy.

OpenStudy (isaiah.feynman):

I rewrite it as \[3\int\limits_{\frac{ \pi }{ 6}}^{\frac{ \pi }{ 2 }} \frac{ \cos(x) }{ \sin(x) } dx\]

OpenStudy (anonymous):

if you substitute sin(x) for x, it should make it a bit easier...

OpenStudy (isaiah.feynman):